OpenStudy (anonymous):

x^2-16 over x^2-8x+16

OpenStudy (luigi0210):

\[\frac{x^2-16}{x^2-8x+16}=\frac{(x+4)(x-4)}{(x-6)(x-2)}\] Finish it from here?

OpenStudy (luigi0210):

*(x-4)(x-4) For the bottom sorry

OpenStudy (anonymous):

so its -1

OpenStudy (luigi0210):

Not exactly: \[\frac{(x+4)(x-4)}{(x-4)(x-4)}=\frac{x+4}{x-4}\]

OpenStudy (anonymous):

\[\frac{ x^2-16 }{ x^2-8x+16 } = \frac{ (x + 4)(x-4) }{ (x - 4)(x-4) }\]

OpenStudy (anonymous):

so a (x-4) from the numerator and a (x-4) from the denominator "cancel out".

OpenStudy (luigi0210):

^

OpenStudy (anonymous):

oh okay , but if it was -4 over 4 it would be -1 right?

OpenStudy (anonymous):

\[\frac{ x^2-16 }{ x^2-8x+16 } = \frac{ (x + 4)(x-4) }{ (x - 4)(x-4) } = \frac{ (x+4) }{ (x-4) }\]

OpenStudy (luigi0210):

Nope, those can't cancel.. you have to leave them as they are

OpenStudy (anonymous):

okay thank you (:

OpenStudy (luigi0210):

Anytime :)